OpenStudy (anonymous):

please tell me this question it's so easy but i don't know why i'm not getting answer 9-8+48 divided by -3/4-8 as show below

OpenStudy (anonymous):



OpenStudy (ranga):

Is the numerator 9-8+48 or 9-80+48 ? In the original posting you have the first one and in the reply box you have the second one.

OpenStudy (anonymous):

sorry 80 is as shown above in DRAW

OpenStudy (ranga):

Numerator: 9 - 80 + 48 = -23 Denominator: -3/4 - 8 = -3/4 - 32/4 = (-3 - 32)/4 = -35/4 Divide: -23 / (-35/4) = 23 * 4 / 35 = 92/35

OpenStudy (wolf1728):

ziqbal103 Since you think you are doing this wrong, you must have the answer of what it should be, correct?